OpenStudy (meilendurcer):

Over the course of the last few years, Lisa has made six donations to her public radio station. If the first of the donations was $7.50, and if each subsequent donation was for 4 times the amount of the previous donation, how much has Lisa donated to her public radio station over the last few years?

OpenStudy (mathstudent55):

Use $7.50 as the first term in a sequence. Then keep multiplying it by 4 until you have 6 terms in the sequence. Then add all 6 terms up.

OpenStudy (meilendurcer):

@mathstudent55 i tried

OpenStudy (meilendurcer):

@mathstudent55 i think i did it wrong though because all i got was 40,000

OpenStudy (mathstudent55):

7.50 4 * 7.50 = 30 30 * 4 = 120 120 * 4 = 480 480 * 4 = 1920 1920 * 4 = 7680 7680 * 4 = 30,720 7.5 + 30 + 120 + 480 + 1920 + 7680 + 30,720 = 40,957.50

OpenStudy (meilendurcer):

yea i got that @mathstudent55

OpenStudy (mathstudent55):

You're correct, then.

OpenStudy (meilendurcer):

@mathstudent55 ok thanks